a bullet of mass m moving horizontally with velocity v hits a block of wood of mass M resting on a smooth horizontal plane. The fraction of energy of the bullet dissipated in the collision itself is?

Grade:12

1 Answers

Abhishek Singh
9 years ago
after collision velocity of syst=v. = mv/(m+M)
initial energy=1/2mv2
Final energy= ½ (M+m)v’2
 
Energy lost = final – initial = ½ mv(M/M+m)
fractional energy lost= energy lost/initial energy= M/M+m